# Discovering Large-Scale Structure at in the C3VO Survey

### v1

@@ 3.3 Detecting Coherent Structure
We obtain the final number of peaks by counting how many peaks have at least five associated points and perform a parametric fit characterized by an identical number of Gaussians to the linked chain (Fig. 3). If the linked chain happens to yield no peaks with at least five points, we simply fit a single Gaussian to the entire linked chain. The initial fitting coefficients for each Gaussian are based on the parameters of the peak, where the mean and amplitude are respectively adopted from the redshift and flux value of the highest flux point in the peak, and initial guess for the redshift dispersion is set by the total redshift range of the points associated to the peak. If no peaks are identified due to an insufficient number of associated points, the initial fitting coefficients for the mean and amplitude of the Gaussian are set by the mean and highest flux value, respectively, of the data points in the linked chain, and we assume a redshift dispersion of 0.01. In order to avoid any likely spurious detections, we only keep the peaks where the converged Gaussian fit satisfies the following conditions: (1) an amplitude not more than 120% of the highest flux value in the peak, (2) a mean offset of from the median redshift of all flux points, and (3) a redshift dispersion of .
@@
We generate the final list of potential overdensities, referred to hereafter as “protostructure candidates”, by sorting the peaks per field from high to low Gaussian amplitude and subsuming together any candidates that fall within a transverse distance of 2 Mpc and . In the case of protostructure candidates that result from two or more subsumed components, the amplitudes of the constituent protostructure candidates from the individual fits are added together in order to determine the Gaussian amplitude of each aggregate candidate, and an amplitude-weighted average of the individual positions and redshifts is additionally calculated to determine the location of the aggregate candidate. Entries in our catalog that are composed of two or more subsumed candidates make up about 10% of our total catalog.
@@ 4 Tests with Simulated Data
The performance of our structure search algorithm is heavily dependent on the choice of SExtractor parameters as well as the linking radius. Less restrictive parameters will yield more protostructure candidates, but will result in large number of spurious candidates. To test the efficacy of our search algorithm, we perform mock observations of simulated data and apply the same density mapping and structure search algorithm to the mock data for which we have absolute knowledge of the structure locations. As we describe in the following sections, these mock observations are designed in such a way as to mimic the quality and approach of the actual observations employed in this study. In this way, in principle, subject to the limitations of the simulations, semi-analytics, and our ability to impose all nuances of our own data on the mock observations, we are able to determine the ability of our data, mapping, and search algorithm to detect structures of differing properties in the early universe.
@@ 4.1 Construction of the Lightcone
For our study, we use the predictions of the GAlaxy Evolution and Assembly (GAEA; De Lucia et al., 2024) semi-analytic model (SAM) to generate simulated galaxy catalogues. The GAEA model, using the version described in Xie et al. (2017), was applied to the dark matter (DM) merger trees of the Millennium Simulation (Springel et al., 2005). This DM simulation is based on particles, each with a mass of and distributed within a box of size comoving Mpc on a side. The simulation is run assuming a CDM model with , , , , and .
Following Zoldan et al. (2017), we generated a lightcone from the output of the GAEA SAM. This lightcone mimics a circular region of the sky with a radius of 2.3 degrees. The lightcone includes galaxies down to a stellar mass of , though is complete down to masses of . The lightcone ranges from to and includes a number of physical and observable properties for model galaxies, such as their RA, Dec, redshift, and membership in DM halos and subhalos. For our study, we make use of the following quantities: RA and Dec, observed redshift (i.e., the cosmological redshift plus the component of the peculiar velocity along the line of sight), and the dust-attenuated observed flux in IRAC1. We also make use of the known membership of each galaxy to its parent halo. In the Millennium DM Simulation, halos are identified by a friends-of-friends technique applied to the dark matter particles, using a linking length of 0.2 in units of the mean interparticle separation. For each halo, the virial mass is also provided as an output of the simulation.
Note that the area of the lightcone is larger than the DM simulation. As such, the lightcone is constructed of a patchwork of simulation cubes taken at single snapshots with different orientations at varying redshifts in order to populate the entire redshift range of the lightcone. Because of the finite size of the cubes, there may be issues where structure is cut off at the interface between two pasted cubes, which results in the possibility of massive systems with only a few sub-halos as members. We identify and eliminate these systems from consideration in our analysis as described in §4.3 and Appendix A.
@@ 4.2 Selecting Lightcone Regions and Mock Observations
For this study, we selected five different, non-overlapping 1x1 deg2 fields in the lightcone separated in RA/Dec. Figure 4 shows the location of each of these regions in the lightcone. A two-dimensional histogram of the masses and redshifts of the protostructures555Here we use the term “protostructure” to indicate the ensemble of halos at a given redshift that comprise a single structure in the simulation.. in these regions is shown in Figure 5. Once these regions were selected, we then assign photometric and spectroscopic redshifts to the simulated galaxies in a way that mimicked the properties of the real data, a process we refer to as “mocking up” the lightcone. For the construction of this catalog, the COSMOS field is chosen as the reference field because it is the largest of the three C3VO fields and contains the majority of our data. We will discuss the mocking up of the other two C3VO fields (ECDFS and CFHTLS-D1) later in this section.
@@ 4.2.1 Construction of the Mock Galaxy Catalogs
For each region selected in the lightcone, we constructed mock galaxy catalogs using the following prescription. We began with the combined photometric and spectroscopic catalog in the COSMOS field, which are described in §2.1.1 and §2.1.2. In order to impose conformity on the number counts of simulated and observed objects, the simulated and observational catalogs were cut at IRAC124.8. At fainter magnitudes, the number counts of simulated objects over the redshift range and appeared to considerably outpace those in the observed data, which likely indicated that the data were not complete to magnitudes fainter than this limit. Objects fainter than IRAC1 were not considered further in any part of our analysis.
@@
In each redshift and magnitude bin we calculate the total number of galaxies with secure spectroscopic redshifts divided by the total number of objects in that bin, where the number of objects in the denominator is set by magnitude and photometric redshift cuts. The redshift bins are set to be wide enough to minimize the effects of scatter across bins due to photometric redshift errors. This quotient sets the spectroscopic redshift fraction for that magnitude and redshift bin. This spectroscopic redshift fraction (SF) for the COSMOS field across all redshift and magnitude bins is referred to hereafter as SF1.0 and sets our fiducial mock spectroscopic observations for the mock catalogs. By definition, SF1.0 carries with it the a posteriori spectroscopic selection function of the COSMOS field, which is then applied to all mocks.
Per redshift and magnitude bin, the spectroscopic redshift fraction in COSMOS ranged from 0.3 for brighter objects (IRAC1) at lower redshift () to 0.001 for the faintest objects (IRAC1 ) at the highest redshifts (). Overall, SF1.0 corresponds to a spectroscopic redshift fraction of 0.07 for a sample cut at IRAC124.8. The equivalent numbers for the ECDFS and CFHTLS-D1 fields were 0.2 and 0.06, respectively, though with the latter cut to a brightness limit of rather than IRAC124.8 due to the properties of the photometric data in that field (see §2.1.1).
In order to approximate the properties of the photometric redshifts in the real data, objects were broken into the same redshift and magnitude bins as were used for the SF calculation. For each magnitude and redshift bin, three quantities were calculated by comparing spectroscopic and photometric redshifts: the bias, the (see Lemaux et al. 2022 and references therein), and the median photometric redshift uncertainty as calculated from the full probability density function (PDF). This latter quantity was denoted . In bins that contained 10 secure spectroscopic redshifts, was adopted as the photometric redshift scatter, otherwise was adopted.
SF | Spectroscopic Fraction |
@@
Note. — SF1.0 is explicitly defined as equal to the spectroscopic redshift fraction for the data in the COSMOS field that is adopted in this study. While the spectroscopic fractions listed here are for all objects brighter than IRAC1, the full spectroscopic selection function is applied in finer magnitude and redshift bins for each SF. Refer to §4.2.2 for more details.
@@ 4.2.2 Mock Map Regions
We chose five 1x1 deg2 regions in the lightcone to perform mock observations (Fig. 4). For a mock observation of a given region of the lightcone, referred to hereafter as “mock X”, where X indicates the region of the lightcone as defined in Figure 4, we generated a spectroscopic catalog using the SF1.0 selection function as a function of magnitude and redshift as determined from the observed data. Spectroscopic redshifts were assigned randomly to galaxies in a given magnitude/redshift bin, and spectroscopic quality flags were assigned in a manner that mimicked the distribution in the real data. For each mock region other than mock 1, we generated only a single spectroscopic redshift catalog using the SF1.0 selection function.
@@
In addition, all objects, regardless of whether they were assigned a spectroscopic redshift, were assigned a photometric redshift using the prescription: , where is the spectroscopic to photometric redshift bias in each magnitude and redshift bin, is a value sampled from a normalized Gaussian distribution for each object, and is either set to or depending on the number of secure spectroscopic redshifts in a given bin. Effective 1 photometric redshift uncertainties for each simulated galaxy were determined by sampling from the PDF of the fractional photometric redshift error (i.e., -)/ and -)/) constructed from the statistics of objects in each magnitude bin in the real data. At the end of this process each simulated galaxy had a photometric redshift with an associated 1 uncertainty. Additionally, those simulated galaxies selected for mock spectral observations were assigned a spectroscopic quality flag and a spectral redshift. These mock galaxy catalogs were then run through the same VMC process as was used for the real data (see §3.2) to generate density reconstructions of all mock realizations of each field. Figure 6 shows an example of region in and around a massive protocluster () at located in the GAEA lightcone and the mock spectroscopic observations for five SF variants. Figure 7 shows the VMC reconstruction of the overdensity field in the same regions for the two extrema of the SF variants, SF0.5 and SF4.0.
@@ 4.3 Obtaining the Final List of Subsumed Structure Candidates
The lightcone is populated with structures scattered across its entire redshift extent. The spacing between structures, however, can vary, such that two structures may be so close together that we would be unable to distinguish them as separate candidates in our VMC maps. We therefore want to construct a structure catalog that encodes this confusion limit in some way so that we can better match the detections in our mock observations and have a clearer measure of the masses of the structures we detect. We did so in the following manner. Using the highest mass structure in the lightcone as a seed, we subsumed all structures that fell within and a transverse separation of 2 Mpc of this structure. These values were chose to roughly match the confusion limit of structures present in our VMC maps. Once all nearby structures were subsumed, their masses are combined into a single protostructure with RA/Dec/z coordinates dictated by a mass-weighted average of the barycenters of all constituent structures. This process continued from the highest mass structure to the lowest mass structure in the lightcone until all single or composite structures in the catalog had no companions within the above limits.
Additionally, we want to account for any artificially underrich structures that may have a confounding effect on measurements of purity, completeness, and any mass-richness relation we attempt. As mentioned in §4.1, such structures likely exist in the lightcone due to structures being artificially cutoff at the interface between two pasted cubes.
We identify these structures by fitting a relation between their masses and numbers of members, divided into redshift bins of . Because we want to take care to not excise systems that are underrich at our adopted magnitude limit of IRAC124.8 due to astrophysical reasons, e.g., a shifting of the luminosity function to fainter magnitudes for a given system resulting in a preponderance of member galaxies that fall below our magnitude cut, we perform this exercise with a simulated galaxy catalog cut at IRAC125.5. For each redshift bin, we fit a linear relation to the logarithmic function of both parameters, iteratively clipping out outliers that are lower in number of members than our best-fit relation. Our underrichness criterion is defined by this same threshold (see Appendix A).
@@ 4.4 Recovery of Simulated Structures
The VMC maps generated from the mock data were then fed into SExtractor in a manner identical to that used for the VMC maps generated from the real data (see §3.3) with two exceptions. Because the mock observations lack the somewhat uneven sampling that is present in the real spectroscopic and photometric data, edge effects, while still present, are less of a concern. As such, we simply mask the bottom 15% of density values in each slice of the mock VMC maps, as opposed to the outer 20-30% area of each slice used for the VMC maps generated from the real data. This choice was made to maximize the area that we can recover signal on. Additionally, the median density of the mocks differs somewhat from that of the real data due to higher order differences between the data and the mock catalogs (Fig. 1). Since the VMC maps generated from the mock data generally exhibit higher average densities than that generated from the real data, we adopt a DETECT_THRESH value of 2.5 for the mock data, which roughly corresponds to 3 values used for the real data.
@@
As we did for the real data, we then generate a list of protostructure candidates from linked detections across different redshift slices in the VMC maps of the mock galaxy catalogs. These candidates then undergo an identical subsuming process as that described in §3.3. In order to test the efficacy of our method both in detecting subsumed structure at different masses and in recovering a signal that is commensurate in strength with the mass of each subsumed structure, we need to create a formalism to match the protostructure candidates we detect to the lightcone structures. This formalism was generated in the following manner.
@@
We have full knowledge of the extent of the member galaxies in RA, Dec and redshift space for each subsumed structure in the lightcone. We match a subsumed structure to a candidate (i.e., consider that simulated structure recovered) if the candidate’s barycenter falls within the minimum and maximum ranges of the rectangular prism of the subsumed structure’s RA and Dec and redshift extents. If the candidate falls within range of multiple subsumed structures, we match the candidate to the subsumed structure that has the closest barycenter in RA and Dec and redshift. Note that this allows for the possibility of more than one candidate within range matching to a given subsumed structure, which may be signal from an individual substructure or simply spurious detections. Typically, however, the Gaussian amplitude of one candidate will dwarf any other matches. In the cases where we have multiple candidates matched to a given subsumed structure, the Gaussian amplitudes of each candidate are added together to eventually compare to the mass of the subsumed structure to which they are matched (Fig. 8).
@@
At the completion of this process, we have a list of subsumed structures and associated Gaussian amplitudes for those structures that have a matched candidate (or multiple candidates). This catalog is used in the next section to assess the purity and completeness of our method for structures of different masses using differing levels of spectroscopic completeness. The catalog is also used in §4.5 to set the rough mass scaling for our candidates.

### v2

@@ 3.3 Detecting Coherent Structure
We obtain the final number of peaks by counting how many peaks have at least five associated points and perform a parametric fit characterized by an identical number of Gaussians to the linked chain (Fig. 3). If the linked chain happens to yield no peaks with at least five points, we simply fit a single Gaussian to the entire linked chain. The initial fitting coefficients for each Gaussian are based on the parameters of the peak, where the mean and amplitude are, respectively, adopted from the redshift and flux value of the highest flux point in the peak, and initial guess for the redshift dispersion is set by the total redshift range of the points associated to the peak. If no peaks are identified due to an insufficient number of associated points, the initial fitting coefficients for the mean and amplitude of the Gaussian are set by the mean and highest flux value, respectively, of the data points in the linked chain, and we assume a redshift dispersion of 0.01. In order to avoid any likely spurious detections, we only keep the peaks where the converged Gaussian fit satisfies the following conditions: (1) an amplitude not more than 120% of the highest flux value in the peak, (2) a mean offset of from the median redshift of all flux points, and (3) a redshift dispersion of .
@@
We generate the final list of potential overdensities, referred to hereafter as “protostructure candidates,” by sorting the peaks per field from high to low Gaussian amplitude and subsuming together any candidates that fall within a transverse distance of 2 Mpc and . In the case of protostructure candidates that result from two or more subsumed components, the amplitudes of the constituent protostructure candidates from the individual fits are added together in order to determine the Gaussian amplitude of each aggregate candidate, and an amplitude-weighted average of the individual positions and redshifts is additionally calculated to determine the location of the aggregate candidate. Entries in our catalog that are composed of two or more subsumed candidates make up about 10% of our total catalog.
The choice of the various parameters described above will affect the completeness and purity of protostructure candidates we find. Less restrictive parameters will yield more protostructure candidates, but will result in a large number of spurious candidates. Ideally, we would like to find the set of values that would maximize both our completeness and purity, as was explored in depth in Hung et al. (2020). However, the parameter space is enormous and would be time-consuming to investigate completely. More care would be necessary for studies that rely on accurate structure number densities, such as cluster mass functions (e.g., Hung et al., 2021). As we are only interested in a general structure search in this work, we adopt values in line with what we previously found to be optimal in Hung et al. (2020) and Hung et al. (2021), finding negligible differences in the narrow parameter space we investigated in this work.
@@ 4 Tests with Simulated Data
To test the efficacy of our search algorithm, we perform mock observations of simulated data and apply the same density mapping and structure search algorithm to the mock data for which we have absolute knowledge of the structure locations. As we describe in the following Sections, these mock observations are designed in such a way as to mimic the quality and approach of the actual observations employed in this study. In this way, in principle, subject to the limitations of the simulations, semianalytics, and our ability to impose all nuances of our own data on the mock observations, we are able to determine the ability of our data, mapping, and search algorithm to detect structures of differing properties in the early Universe.
@@ 4.1 Construction of the Lightcone
For our study, we use the predictions of the GAEA (De Lucia et al., 2024) semianalytic model (SAM) to generate simulated galaxy catalogs. The GAEA model, using the version described in Xie et al. (2017), was applied to the dark matter (DM) merger trees of the Millennium Simulation (Springel et al., 2005). This DM simulation is based on particles, each with a mass of and distributed within a box of size comoving Mpc on a side. The simulation is run on a slightly different cosmology from the rest of this work, assuming a CDM model with , , , , and .
Following Zoldan et al. (2017), we generated a lightcone from the output of the GAEA SAM. This lightcone mimics a circular region of the sky with a radius of 2.3 degrees. The lightcone includes galaxies down to a stellar mass of , though is complete down to masses of . The lightcone ranges from to and includes a number of physical and observable properties for model galaxies, such as their R.A., decl., redshift, and membership in DM halos and subhalos. For our study, we make use of the following quantities: R.A. and decl., observed redshift (i.e., the cosmological redshift plus the component of the peculiar velocity along the line-of-sight), and the dust-attenuated observed flux in IRAC1. We also make use of the known membership of each galaxy to its parent halo. In the Millennium DM Simulation, halos are identified by a friends-of-friends technique applied to the dark matter particles, using a linking length of 0.2 in units of the mean interparticle separation. For each halo, the virial mass is also provided as an output of the simulation.
Note that the area of the lightcone is larger than the DM simulation. As such, the lightcone is constructed of a patchwork of simulation cubes taken at single snapshots with different orientations at varying redshifts in order to populate the entire redshift range of the lightcone. Because of the finite size of the cubes, there may be issues where structure is cut off at the interface between two pasted cubes, which results in the possibility of massive systems with only a few subhalos as members. We identify and eliminate these systems from consideration in our analysis as described in Section 4.3 and Appendix A.
@@ 4.2 Selecting Lightcone Regions and Mock Observations
For this study, we selected five different, nonoverlapping 11 deg2 fields in the lightcone separated in R.A./decl. Figure 4 shows the location of each of these regions in the lightcone. A two-dimensional histogram of the masses and redshifts of the protostructures555Here we use the term “protostructure” to indicate the ensemble of halos at a given redshift that comprise a single structure in the simulation.. in these regions is shown in Figure 5. Once these regions were selected, we then assign photometric and spectroscopic redshifts to the simulated galaxies in a way that mimicked the properties of the real data, a process we refer to as “mocking up” the lightcone, which we describe in the following subsection. For the construction of this catalog, the COSMOS field is chosen as the reference field, because it is the largest of the three C3VO fields and contains the majority of our data. We will discuss the mocking up of the other two C3VO fields (ECDFS and CFHTLS-D1) later in this Section.
@@ 4.2.1 Construction of the Mock Galaxy Catalogs
For each region selected in the lightcone, we constructed mock galaxy catalogs using the following prescription. We began with the combined photometric and spectroscopic catalog in the COSMOS field, which are described in Sections 2.1.1 and 2.1.2. In order to impose conformity on the number counts of simulated and observed objects, the simulated and observational catalogs were cut at IRAC124.8. At fainter magnitudes, the number counts of simulated objects over the redshift range and appeared to considerably outpace those in the observed data, which likely indicated that the data were not complete to magnitudes fainter than this limit. Objects fainter than IRAC1 were not considered further in any part of our analysis.
@@
In each redshift and magnitude bin we calculate the total number of galaxies with secure spectroscopic redshifts divided by the total number of objects in that bin, where the number of objects in the denominator is set by magnitude and photometric redshift cuts. The redshift bins are set to be wide enough to minimize the effects of scatter across bins due to photometric redshift errors. This quotient sets the global spectroscopic redshift fraction for that magnitude and redshift bin. The spectroscopic redshift fraction (SF) for the COSMOS field across all redshift and magnitude bins is referred to hereafter as SF1.0 and sets our fiducial mock spectroscopic observations for the mock catalogs. By definition, SF1.0 carries with it the a posteriori spectroscopic selection function of the COSMOS field, which is then applied to all mocks.
The global spectroscopic redshift fraction in COSMOS per redshift and magnitude bin ranged from 0.3 for brighter objects (IRAC1) at lower redshift () to 0.001 for the faintest objects (IRAC1 ) at the highest redshifts (). Overall, SF1.0 corresponds to a global spectroscopic redshift fraction of 0.07 for a sample cut at IRAC124.8. The equivalent numbers for the ECDFS and CFHTLS-D1 fields were 0.2 and 0.06, respectively, though with the latter cut to a brightness limit of rather than IRAC124.8 due to the properties of the photometric data in that field (see Section 2.1.1).
In order to approximate the properties of the photometric redshifts in the real data, objects were broken into the same redshift and magnitude bins as were used for the SF calculation. For each magnitude and redshift bin, three quantities were calculated by comparing spectroscopic and photometric redshifts: the bias, the (see Lemaux et al. 2022 and references therein), and the median photometric redshift uncertainty as calculated from the full PDF. This latter quantity was denoted . In bins that contained 10 secure spectroscopic redshifts, was adopted as the photometric redshift scatter; otherwise was adopted.
SF | Global Spectroscopic Fraction |
@@
Note. — SF1.0 is explicitly defined as equal to the global spectroscopic redshift fraction for the data in the COSMOS field that is adopted in this study. While the spectroscopic fractions listed here are for all objects brighter than IRAC1, the full spectroscopic selection function is applied in finer magnitude and redshift bins for each SF. Refer to Section 4.2.2 for more details.
@@ 4.2.2 Mock Map Regions
We chose five 11 deg2 regions in the lightcone to perform mock observations (Fig. 4). For a mock observation of a given region of the lightcone, referred to hereafter as “mock X,” where X indicates the region of the lightcone as defined in Figure 4, we generated a spectroscopic catalog using the SF1.0 selection function as a function of magnitude and redshift as determined from the observed data. Spectroscopic redshifts were assigned randomly to galaxies in a given magnitude/redshift bin, and spectroscopic quality flags were assigned in a manner that mimicked the distribution in the real data. For each mock region other than mock 1, we generated only a single spectroscopic redshift catalog using the SF1.0 selection function.
@@
In addition, all objects, regardless of whether they were assigned a spectroscopic redshift, were assigned a photometric redshift using the prescription: , where is the spectroscopic to photometric redshift bias in each magnitude and redshift bin, is a value sampled from a normalized Gaussian distribution for each object, and is either set to or depending on the number of secure spectroscopic redshifts in a given bin. Effective 1 photometric redshift uncertainties for each simulated galaxy were determined by sampling from the PDF of the fractional photometric redshift error (i.e., -)/ and -)/) constructed from the statistics of objects in each magnitude bin in the real data. At the end of this process, each simulated galaxy had a photometric redshift with an associated 1 uncertainty. Additionally, those simulated galaxies selected for mock spectral observations were assigned a spectroscopic quality flag and a spectral redshift. These mock galaxy catalogs were then run through the same VMC process as was used for the real data (see Section 3.2) to generate density reconstructions of all mock realizations of each field. Figure 6 shows an example of the region in and around a massive protocluster () at located in the GAEA lightcone and the mock spectroscopic observations for five SF variants. Figure 7 shows the VMC reconstruction of the overdensity field in the same regions for the two extrema of the SF variants, SF0.5 and SF4.0.
@@ 4.3 Obtaining the Final List of Subsumed Structure Candidates
The lightcone is populated with structures scattered across its entire redshift extent. The spacing between structures, however, can vary, such that two structures may be so close together that we would be unable to distinguish them as separate candidates in our VMC maps. We therefore want to construct a structure catalog that encodes this confusion limit in some way so that we can better match the detections in our mock observations and have a clearer measure of the masses of the structures we detect. We did so in the following manner. Using the highest mass structure in the lightcone as a seed, we subsumed all structures that fell within and a transverse separation of 2 Mpc of this structure. These values were chose to roughly match the confusion limit of structures present in our VMC maps. Once all nearby structures were subsumed, their masses are combined into a single protostructure with R.A./decl./ coordinates dictated by a mass-weighted average of the barycenters of all constituent structures. This process continued from the highest mass structure to the lowest mass structure in the lightcone until all single or composite structures in the catalog had no companions within the above limits.
Additionally, we want to account for any artificially under-rich structures that may have a confounding effect on measurements of purity, completeness, and any mass–richness relation we attempt. As mentioned in Section 4.1, such structures likely exist in the lightcone due to structures being artificially cut off at the interface between two pasted cubes.
We identify these structures by fitting a relation between their masses and numbers of members, divided into redshift bins of . Because we want to take care to not excise systems that are under-rich at our adopted magnitude limit of IRAC124.8 due to astrophysical reasons (e.g., a shifting of the luminosity function to fainter magnitudes for a given system resulting in a preponderance of member galaxies that fall below our magnitude cut), we perform this exercise with a simulated galaxy catalog cut at IRAC125.5. For each redshift bin, we fit a linear relation to the logarithmic function of both parameters, iteratively clipping out outliers that are lower in number of members than our best-fit relation. Our under-richness criterion is defined by this same threshold (see Appendix A).
@@ 4.4 Recovery of Simulated Structures
The VMC maps generated from the mock data were then fed into SExtractor in a manner identical to that used for the VMC maps generated from the real data (see Section 3.3) with two exceptions. Because the mock observations lack the somewhat uneven sampling that is present in the real spectroscopic and photometric data, edge effects, while still present, are less of a concern. As such, we simply mask the bottom 15% of density values in each slice of the mock VMC maps, as opposed to the outer 20–30% area of each slice used for the VMC maps generated from the real data. This choice was made to maximize the area that we can recover a signal on. Additionally, the median density of the mocks differs somewhat from that of the real data due to higher-order differences between the data and the mock catalogs (Fig. 1). Since the VMC maps generated from the mock data generally exhibit higher average densities than that generated from the real data, we adopt a DETECT_THRESH value of 2.5 for the mock data, which roughly corresponds to 3 values used for the real data (see Fig. 1).
@@
As we did for the real data, we then generate a list of protostructure candidates from linked detections across different redshift slices in the VMC maps of the mock galaxy catalogs. These candidates then undergo an identical subsuming process to that described in Section 3.3. In order to test the efficacy of our method both in detecting subsumed structure at different masses and in recovering a signal that is commensurate in strength with the mass of each subsumed structure, we need to create a formalism to match the protostructure candidates we detect to the lightcone structures. This formalism was generated in the following manner.
@@
We have full knowledge of the extent of the member galaxies in R.A., decl., and redshift space for each subsumed structure in the lightcone. We match a subsumed structure to a candidate (i.e., consider that simulated structure recovered) if the candidate’s barycenter falls within the minimum and maximum ranges of the rectangular prism of the subsumed structure’s R.A. and decl. and redshift extents. If the candidate falls within range of multiple subsumed structures, we match the candidate to the subsumed structure that has the closest barycenter in R.A. and decl. and redshift. Note that this allows for the possibility of more than one candidate within range matching to a given subsumed structure, which may be a signal from an individual substructure or simply spurious detections. Typically, however, the Gaussian amplitude of one candidate will dwarf any other matches. In the cases where we have multiple candidates matched to a given subsumed structure, the Gaussian amplitudes of each candidate are added together to eventually compare to the mass of the subsumed structure to which they are matched (Fig. 8).
@@
At the completion of this process, we have a list of subsumed structures and associated Gaussian amplitudes for those structures that have a matched candidate (or multiple candidates). This catalog is used in the next Section to assess the purity and completeness of our method for structures of different masses using differing levels of spectroscopic completeness. The catalog is also used in Section 4.5 to set the rough mass scaling for our candidates.